# Anisotropic linear waves and breakdown of the momentum expansion in spin magnetohydrodynamics

### v1

@@ Abstract
We formulate spin magnetohydrodynamics (MHD) by including the magnetic-flux and total angular momentum conservation laws in the hydrodynamic framework. To specify the local angular momentum conservation, we choose the totally antisymmetric spin current. The entropy-current analysis allows for ten dissipative first-order transport coefficients including anisotropic spin relaxation rates and the conversion rate between a vorticity (shear) to a symmetric stress (antisymmetric torque). By employing the linear-mode analysis, we solve the first-order spin MHD equations to determine the dispersion relations with the complete information of anisotropy retained. Our analytic solutions indicate that the small-momentum expansion is spoiled by blow up of the higher-order terms when the angle between the momentum and the magnetic field approaches the right angle. This also reveals the existence of another expansion parameter, and, in light of it, we provide solutions in an alternative series expression beyond the critical angle. We confirm that these two series expansions work well in the appropriate angle ranges as compared with numerical results.
@@ I Introduction
In the realm of nuclear physics, the study of spin polarization in relativistic nucleus-nucleus collisions has unveiled a fascinating exploration domain. High-energy collisions generate vortical fields within the Quark-Gluon Plasma (QGP) through the initial orbital angular momentum, leading to the global spin polarization of $\Lambda$ and $\bar{\Lambda}$ hyperons and the alignment of vector mesons via spin-orbit couplings [1, 2, 3, 4, 5, 6]. The polarization of hyperons was observed in experiment by STAR at RHIC and ALICE at LHC [7, 8, 9, 10, 11], while the alignment of vector mesons was also observed in recent years [12, 13], representing a pivotal moment in understanding the QCD matter under extreme conditions.
@@
Theoretical investigations have also been devoted to developing relativistic frameworks for spin dynamics in QGP. Early theoretical projections suggested that the spin polarization of particles would be affected by the thermal vorticity tensor at the freeze-out surface, intricately associated with the fluid’s angular velocity and acceleration. Various theoretical models are adopted to include the thermal vorticity and have successfully described the global polarization using the modified Cooper-Frye formula [4, 14, 15, 16, 17, 18]. Nevertheless, the experimental results on the local $\Lambda$ spin polarization pose a spin sign puzzle that cannot be solely explained by thermal effects [19, 20, 21], i.e., theoretical calculations give the opposite azimuthal angle dependence of $\Lambda$ polarization to the experimental results. This deviation may be attributed to the non-equilibrium evolution of spin degrees of freedom, inspiring theoretical research on hydrodynamic descriptions that include spin effects. Substantial theoretical studies have been conducted on the relativistic spin hydrodynamics based on the entropy analysis or the second law of thermodynamics [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], quantum kinetic theories for fermions [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54], holographic approaches [55, 56, 57], effective action approach [58, 59], and the statistical density operator methods [60, 61, 62, 63, 26, 27, 64]. The spin relaxation rate was also evaluated with the Kubo formula [48]. Recently, the shear-induced polarization (SIP) [65] has been proposed as a source triggering the local spin polarization. Including them in hydrodynamic simulations seems to qualitatively resolve the sign issue [66, 67] (see also Ref. [68, 69]).
@@
Besides the experimental observation of the spin polarization, strong magnetic fields are expected to be generated in the non-central collisions (see, e.g., Refs. [70, 71, 72] for reviews and references therein). These magnetic fields are thought to be transient as they will fade out rapidly in the heavy-ion collision, but may result in some important phenomena like chiral magnetic effect [73, 74, 75, 76] and chiral magnetic wave [77]. Furthermore, the strong magnetic field may also have a significant impact on the spin polarization. Unlike rotational effects, a magnetic field induces opposite spin polarization of particles and antiparticles as they carry opposite charges. This could account for the observed deviations in the polarization of $\Lambda$ and $\bar{\Lambda}$ in heavy-ion collisions [8]. Recently, magnetic-field effects were discussed as sources of not only the splitting between $\Lambda$ and $\bar{\Lambda}$ [78, 79, 80] but also the local spin polarization [81]. The interplay between the magnetic fields and spin dynamics in the QGP draws considerable interest.
@@
In this study, we explore formulation and linear-mode analysis of the relativistic spin magnetohydrodynamics (MHD), extending our study on the first-order MHD [82]. We plug spin dynamics into relativistic MHD that has been formulated based on the conservation of magnetic flux in recent years [83, 84, 63, 85] (see Ref. [86] for a review). We will find anisotropic rotational viscosities and a new transport coefficient that converts the vorticity to a symmetric stress and the shear to an antisymmetric torque. We also solve the system of hydrodynamic equations for the linear perturbation in the vicinity of an equilibrium configuration. Note that, while the linear-mode analyses have been applied to MHD [83, 87, 85, 82] and spin hydrodynamics [22, 63, 27, 33, 30, 31], the linear waves in the spin MHD has not been addressed, which is exactly what we will present in this work. We also note that spin MHD was recently discussed with the kinetic theory approach [88, 89].
@@
In Sec.II, we provide the derivation details of the hydrodynamic equations for spin MHD. After Sec.III, we focus on the linear-mode analysis near an equilibrium state. We elaborate on the solutions for the linear waves in Sec.IV. In this part, a pair of the Alfven waves and two pairs of the magneto-sonic waves are found in the MHD sector, while three damping modes are found in the spin sector. We provide a remedy for the breakdown of the momentum expansion. A comparison with our previous work on MHD [82] is also given. We give a summary in Sec.V.
@@ II.1 Conservation laws
where $\Theta^{\mu\nu}$, $J^{\mu\alpha\beta}$, ${\tilde{F}}^{\mu\nu}$ are the energy-momentum tensor, the angular momentum tensor, and the dual electromagnetic field strength tensor, respectively.
@@
$e=u_{\mu}u_{\nu}{\Theta^{\mu\nu}},\quad S^{\alpha\beta}=-u_{\mu}\Sigma^{\mu\alpha\beta},\quad B^{\mu}={\color[rgb]{0.7,0.3,0.9}{-}}{\tilde{F}}^{\mu\nu}u_{\nu}.$ (4)
@@ II.2 Entropy-current analysis
$+\partial_{\mu}(s_{(1)}^{\mu}+\beta u_{\nu}\Theta^{\mu\nu}_{(1)}+\beta\omega_{\alpha\beta}\Sigma^{\mu\alpha\beta}_{(1)}-\beta H_{\nu}\tilde{F}^{\mu\nu}_{(1)}),$
@@
$-T\left(\begin{array}[]{ll}\eta^{\mu\nu\rho\sigma}&\xi^{\mu\nu\rho\sigma}\\ \xi^{\prime\mu\nu\rho\sigma}&\gamma^{\mu\nu\rho\sigma}\end{array}\right)\left(\begin{array}[]{l}\partial_{(\rho}\beta_{\sigma)}\\ \partial_{[\rho}\beta_{\sigma]}-2\beta\omega_{\rho\sigma}\end{array}\right).$
@@
where $\zeta_{\times}=\zeta_{\times}^{\prime}$ by virtue of Onsager’s reciprocal relation. These viscous tensors are analogs of those in the presence of a strong vorticity [29].
@@
To ensure Eq. (10) to be satisfied, the general form of the fourth-rank tensor $\rho^{\mu\nu\rho\sigma}$ can be constructed as [83, 90, 86]
@@ III.1 Equations for the energy-momentum tensor
00 0 $=$ $\partial_{0}\delta\epsilon+\frac{B}{\mu_{m}}\partial_{0}\delta B$ (27a)
@@
00 0 $=$ $h{\partial_{0}}{\delta u_{x}}+c_{s}^{2}{\partial_{x}}{\delta\epsilon}-\frac{B}{\mu_{m}}{\partial_{z}}\delta{B_{x}}+\frac{B}{\mu_{m}}\partial_{x}\delta B_{z}$ (27b)
@@
00 0 $=$ $h{\partial_{0}}{\delta u_{y}}+c_{s}^{2}{\partial_{y}}{\delta\epsilon}-\frac{B}{\mu_{m}}{\partial_{z}}\delta{B_{y}}+\frac{B}{\mu_{m}}\partial_{y}\delta B_{z}$ (27c)
@@
00 0 $=$ $c_{s}^{2}{\partial_{z}}\delta\epsilon+(h-\frac{B^{2}}{\mu_{m}}){\partial_{0}}{\delta u_{z}}$ (27d)
@@ III.2 Equations for the magnetic field
00 0 $=$ $\partial_{i}\delta B^{i},$ (31a)
00 0 $=$ $B\partial_{z}\delta u_{x}-\partial_{0}\delta B_{x}$ (31b)
@@
00 0 $=$ $B\partial_{z}\delta u_{y}-\partial_{0}\delta B_{y}$ (31c)
@@
00 0 $=$ $-B\partial_{\perp\mu}\delta u^{\mu}_{\perp}-\partial_{0}\delta B_{z}$ (31d)
@@ III.3 Equations for the spin angular momentum
00 0 $=$ $(\Gamma_{\parallel}+\partial_{0})\delta\sigma_{x}$
@@
00 0 $=$ $(\Gamma_{\parallel}+\partial_{0})\delta\sigma_{y}$
@@
00 0 $=$ $(\Gamma_{\perp}+\partial_{0})\delta\sigma_{z}+2\big{[}\,\gamma_{\perp}(\partial_{x}\delta u_{y}-\partial_{y}\delta u_{x})\,\big{]},$ (35c)
@@ III.4 Summary of linearized equations
00 0 $=$ $h{\partial_{0}}{\delta u_{y}}-\frac{h}{B}v_{A}^{2}{\partial_{z}}\delta{B_{y}}$
@@
00 0 $=$ $B\partial_{z}\delta u_{y}-\partial_{0}\delta B_{y}+\rho^{\prime}_{\perp}\partial_{z}^{2}\delta B_{y}+\rho^{\prime}_{\parallel}\partial_{x}^{2}\delta B_{y},$ (37b)
00 0 $=$ $(\Gamma_{\parallel}+\partial_{0})\delta\sigma_{x}-2(\gamma_{\parallel}-\xi_{\parallel})\partial_{z}\delta u_{y},$ (37c)
00 0 $=$ $(\Gamma_{\perp}+\partial_{0})\delta\sigma_{z}+2\gamma_{\perp}\partial_{x}\delta u_{y}.$ (37d)
@@
00 0 $=$ $\partial_{0}\delta\epsilon+\frac{h}{B}v_{A}^{2}\partial_{0}\delta B+h\partial_{x}\delta u_{x}+h(1-v_{A}^{2})\partial_{z}\delta u_{z},$
00 0 $=$ $h{\partial_{0}}{\delta u_{x}}+c_{s}^{2}{\partial_{x}}{\delta\epsilon}-\frac{h}{B}v_{A}^{2}({\partial_{z}}\delta{B_{x}}-\partial_{x}\delta B_{z})$ (40b)
@@
00 0 $=$ $c_{s}^{2}{\partial_{z}}\delta\epsilon+h(1-v_{A}^{2}){\partial_{0}}{\delta u_{z}}-(\zeta_{\times}+\eta_{\parallel}-\gamma_{\parallel})\partial_{z}\partial_{x}\delta u_{x}$ (40c)
@@
00 0 $=$ $B\partial_{z}\delta u_{x}-\partial_{0}\delta B_{x}$ (40d)
@@
00 0 $=$ $(\Gamma_{\parallel}+\partial_{0})\delta\sigma_{y}$ (40e)

### v2

@@ Abstract
We formulate spin magnetohydrodynamics (MHD) by including the magnetic-flux and total angular momentum conservation laws in the hydrodynamic framework. To specify the local angular momentum conservation, we choose the totally antisymmetric spin current. The entropy-current analysis allows for ten dissipative first-order transport coefficients including anisotropic spin relaxation rates and the conversion rate between a vorticity (shear) to a symmetric stress (antisymmetric torque), as well as anisotropic viscosities and resistivities. By employing the linear-mode analysis, we solve the first-order spin MHD equations to determine the dispersion relations with the complete information of anisotropy retained. Our analytic solutions indicate that the small-momentum expansion is spoiled by blow up of the higher-order terms when the angle between the momentum and the magnetic field approaches the right angle. This also reveals the existence of another expansion parameter, and, in light of it, we provide solutions in an alternative series expression beyond the critical angle. We confirm that these two series expansions work well in the appropriate angle ranges as compared with numerical results. Building on our findings regarding the breakdown of the small-momentum expansion in first-order theory, we proceed to discussing how these first-order solutions are modified when we include the relaxation dynamics for dissipative modes with the Israel-Stewart framework. We find that, due to the presence of the critical behavior in the first-order solutions, there remains a diffusive window even after the relaxation dynamics is introduced.
@@ I Introduction
In the realm of nuclear physics, the study of spin polarization in relativistic nucleus-nucleus collisions has unveiled a fascinating exploration domain. High-energy collisions generate vortical fields within the quark-gluon plasma (QGP) through the initial orbital angular momentum, leading to the global spin polarization of $\Lambda$ and $\bar{\Lambda}$ hyperons and the alignment of vector mesons via spin-orbit couplings [1, 2, 3, 4, 5, 6]. The polarization of hyperons was observed in experiments by STAR at RHIC and ALICE at the LHC [7, 8, 9, 10, 11], while the alignment of vector mesons was also observed in recent years [12, 13], representing a pivotal moment in understanding the QCD matter under extreme conditions.
@@
Theoretical investigations have also been devoted to developing relativistic frameworks for spin dynamics in QGP. Early theoretical projections suggested that the spin polarization of particles would be affected by the thermal vorticity tensor at the freeze-out surface, intricately associated with the fluid’s angular velocity and acceleration. Various theoretical models are adopted to include the thermal vorticity and have successfully described the global polarization using the modified Cooper-Frye formula [4, 14, 15, 16, 17, 18]. Nevertheless, the experimental results on the local $\Lambda$ spin polarization pose a spin sign puzzle that cannot be solely explained by thermal effects [19, 20, 21], i.e., theoretical calculations give the opposite azimuthal angle dependence of $\Lambda$ polarization to the experimental results. This deviation may be attributed to the non-equilibrium evolution of spin degrees of freedom, inspiring theoretical research on hydrodynamic descriptions that include spin effects. Substantial theoretical studies have been conducted on the relativistic spin hydrodynamics based on the entropy analysis or the second law of thermodynamics [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], quantum kinetic theories for fermions [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54], holographic approaches [55, 56, 57], effective action approach [58, 59], and the statistical density operator methods [60, 61, 62, 63, 26, 27, 64]. Different from hydrodynamic modes, spin density is not a strictly conserved charge since the spatial rotational symmetry only guarantees the conservation of the total angular momentum including the orbital part. Therefore, there is not a a priori justification of spin hydrodynamics as a (quasi-)hydrodynamics, which should be instead justified posteriori by evaluating the spin relaxation rate. Evaluations of the spin relaxation rate have been addressed recently in Refs. [63, 48, 49, 65, 66] for relativistic heavy-ion collisions. Also, the shear-induced polarization [67] has been proposed as a source triggering the local spin polarization. Including them in hydrodynamic simulations seems to qualitatively resolve the sign issue [68, 69] (see also Refs. [70, 71]).
@@
Besides the experimental observation of the spin polarization, strong magnetic fields are expected to be generated in the non-central collisions (see, e.g., Refs. [72, 73, 74, 75] for reviews and references therein). These magnetic fields are thought to be transient as they will fade out rapidly in the heavy-ion collision, but may result in some important phenomena like chiral magnetic effect [76, 77, 78, 79] and chiral magnetic wave [80]. Recently, there have been developments in the numerical simulation of relativistic magneto-hydrodynamics(MHD) [81, 82, 83, 84, 85]. Furthermore, the strong magnetic field may also have a significant impact on the spin polarization. Unlike rotational effects, a magnetic field induces opposite spin polarization of particles and antiparticles as they carry opposite charges. Though there is still no experimental certainty, the deviations in the polarization of $\Lambda$ and $\bar{\Lambda}$ in heavy-ion collisions have drawn a lot of attention recently [8], which may be accounted by the effect of the magnetic field. Recently, magnetic-field effects were discussed as sources of not only the splitting between $\Lambda$ and $\bar{\Lambda}$ [86, 87, 88] but also the local spin polarization [89]. The interplay between the magnetic fields and spin dynamics in the QGP draws considerable interest.
@@
In this study, we explore formulation and linear-mode analysis of the relativistic spin MHD, extending our study on the first-order MHD [90]. We plug spin dynamics into relativistic MHD that has been formulated in recent years with anisotropic viscosities and resistivities [91, 92, 93, 75, 94, 95, 63, 96, 97] (see Ref. [98] for a review). We will find anisotropic rotational viscosities and a new transport coefficient that converts the vorticity to a symmetric stress and the shear to an antisymmetric torque. We also solve the system of hydrodynamic equations for the linear perturbation in the vicinity of an equilibrium configuration. Note that, while the linear-mode analyses have been applied to MHD [91, 99, 96, 90] and spin hydrodynamics [22, 63, 27, 33, 30, 31], the linear waves in the spin MHD has not been addressed, which is exactly what we will present in this work. We also note that spin MHD was recently discussed with the kinetic theory approach [100, 101]. In this paper, we also introduce relaxation dynamics with the Israel-Stewart (IS) framework. This offers an application of the IS framework to anisotropic systems. We will find that there remains a diffusive window, of which the boundary is specified by a direction of momentum as well as its magnitude.
@@
In Sec.II, we provide the derivation details of the hydrodynamic equations for spin MHD. In Sec.III, we focus on the linear-mode analysis near an equilibrium state. We elaborate on the solutions for the linear waves in Sec.IV. In this part, a pair of the Alfven waves and two pairs of the magneto-sonic waves are found in the MHD sector, while three damping modes are found in the spin sector. We provide a remedy for the breakdown of the momentum expansion. A comparison with our previous work on MHD [90] is also given. We give a summary in Sec.V.
@@ II.1 Conservation laws
where $\Theta^{\mu\nu}$, $J^{\mu\alpha\beta}$, and ${\tilde{F}}^{\mu\nu}$ are the energy-momentum tensor, the angular momentum tensor, and the dual electromagnetic field strength tensor, respectively.
@@
$e=u_{\mu}u_{\nu}{\Theta^{\mu\nu}},\quad S^{\alpha\beta}=-u_{\mu}\Sigma^{\mu\alpha\beta},\quad B^{\mu}={-}{\tilde{F}}^{\mu\nu}u_{\nu}.$ (4)
@@ II.2 Entropy-current analysis
[…]
@@
$-T\left(\begin{array}[]{l}1\\ 1\\ \end{array}\right)\left(\begin{array}[]{ll}\eta^{\mu\nu\rho\sigma}&\xi^{\mu\nu\rho\sigma}\\ \xi^{\prime\mu\nu\rho\sigma}&\gamma^{\mu\nu\rho\sigma}\end{array}\right)\left(\begin{array}[]{l}\partial_{(\rho}\beta_{\sigma)}\\ \partial_{[\rho}\beta_{\sigma]}-2\beta\omega_{\rho\sigma}\end{array}\right).$
@@
where $\zeta_{\times}=\zeta_{\times}^{\prime}$ by virtue of Onsager’s reciprocal relation. These viscous tensors are analogs of those in the presence of a strong vorticity [29]. The transport coefficients, $\gamma_{\perp},\,\gamma_{\parallel},\,\xi_{\parallel}$, in the antisymmetric part control the spin relaxation rate in Eq. (3). [^1] One can confirm this in the solutions for the linear-mode analysis in a later section [see, e.g., Eqs. (54) and (56)]. Therefore, spin density persists near, albeit not strictly in, an equilibrium and the current framework works as a quasi-hydrodynamics when these transport coefficients are small enough. The symmetric viscous tensor $\eta^{\mu\nu\rho\sigma}$ reduces to the familiar viscous tensor in the isotropic limit $\zeta_{\parallel}=\zeta_{\perp}=\zeta_{\times}$ and $\eta_{\parallel}=\eta_{\perp}$. The antisymmetric viscous tensor $\gamma^{\mu\nu\rho\sigma}$ also reduces to that without a magnetic field [22] in the isotropic limit $\gamma_{\parallel}=\gamma_{\perp}$. The cross term $\xi^{\mu\nu\rho\sigma}$ does not exist without $b^{\mu}$. Therefore, the most general tensor structure in Eq. (14) works for an arbitrary magnitude of a magnetic field.
@@
To ensure Eq. (10) is satisfied, the general form of the fourth-rank tensor $\rho^{\mu\nu\rho\sigma}$ can be constructed as [91, 102, 98]
@@ III.1 Equations for the energy-momentum tensor
$0$ $=$ $\partial_{0}\delta\epsilon+\frac{B}{\mu_{m}}\partial_{0}\delta B$ (29a)
@@
$0$ $=$ $h{\partial_{0}}{\delta u_{x}}+c_{s}^{2}{\partial_{x}}{\delta\epsilon}-\frac{B}{\mu_{m}}{\partial_{z}}\delta{B_{x}}+\frac{B}{\mu_{m}}\partial_{x}\delta B_{z}$ (29b)
@@
$0$ $=$ $h{\partial_{0}}{\delta u_{y}}+c_{s}^{2}{\partial_{y}}{\delta\epsilon}-\frac{B}{\mu_{m}}{\partial_{z}}\delta{B_{y}}+\frac{B}{\mu_{m}}\partial_{y}\delta B_{z}$ (29c)
@@
$0$ $=$ $c_{s}^{2}{\partial_{z}}\delta\epsilon+(h-\frac{B^{2}}{\mu_{m}}){\partial_{0}}{\delta u_{z}}$ (29d)
@@ III.2 Equations for the magnetic field
$0$ $=$ $\partial_{i}\delta B^{i},$ (33a)
$0$ $=$ $B\partial_{z}\delta u_{x}-\partial_{0}\delta B_{x}$ (33b)
@@
$0$ $=$ $B\partial_{z}\delta u_{y}-\partial_{0}\delta B_{y}$ (33c)
@@
$0$ $=$ $-B\partial_{\perp\mu}\delta u^{\mu}_{\perp}-\partial_{0}\delta B_{z}$ (33d)
@@ III.3 Equations for the spin angular momentum
$0$ $=$ $(\Gamma_{\parallel}+\partial_{0})\delta\sigma_{x}$
@@
$0$ $=$ $(\Gamma_{\parallel}+\partial_{0})\delta\sigma_{y}$
@@
$0$ $=$ $(\Gamma_{\perp}+\partial_{0})\delta\sigma_{z}+2\big[\,\gamma_{\perp}(\partial_{x}\delta u_{y}-\partial_{y}\delta u_{x})\,\big],$ (37c)
@@ III.4 Summary of linearized equations
$0$ $=$ $h{\partial_{0}}{\delta u_{y}}-\frac{h}{B}v_{A}^{2}{\partial_{z}}\delta{B_{y}}$
@@
$0$ $=$ $B\partial_{z}\delta u_{y}-\partial_{0}\delta B_{y}+\rho^{\prime}_{\perp}\partial_{z}^{2}\delta B_{y}+\rho^{\prime}_{\parallel}\partial_{x}^{2}\delta B_{y},$ (39b)
$0$ $=$ $(\Gamma_{\parallel}+\partial_{0})\delta\sigma_{x}-2(\gamma_{\parallel}-\xi_{\parallel})\partial_{z}\delta u_{y},$ (39c)
$0$ $=$ $(\Gamma_{\perp}+\partial_{0})\delta\sigma_{z}+2\gamma_{\perp}\partial_{x}\delta u_{y}.$ (39d)
@@
$0$ $=$ $\partial_{0}\delta\epsilon+\frac{h}{B}v_{A}^{2}\partial_{0}\delta B+h\partial_{x}\delta u_{x}+h(1-v_{A}^{2})\partial_{z}\delta u_{z},$
$0$ $=$ $h{\partial_{0}}{\delta u_{x}}+c_{s}^{2}{\partial_{x}}{\delta\epsilon}-\frac{h}{B}v_{A}^{2}({\partial_{z}}\delta{B_{x}}-\partial_{x}\delta B_{z})$ (42b)
@@
$0$ $=$ $c_{s}^{2}{\partial_{z}}\delta\epsilon+h(1-v_{A}^{2}){\partial_{0}}{\delta u_{z}}-(\zeta_{\times}+\eta_{\parallel}-\gamma_{\parallel})\partial_{z}\partial_{x}\delta u_{x}$ (42c)
@@
$0$ $=$ $B\partial_{z}\delta u_{x}-\partial_{0}\delta B_{x}$ (42d)
@@
$0$ $=$ $(\Gamma_{\parallel}+\partial_{0})\delta\sigma_{y}$ (42e)
@@
Corresponding to the number of above equations, we will find nine modes. While six of them are modifications of the conventional gapless MHD modes, the remaining three are gapped spin modes in the totally antisymmetric pseudo-gauge.